# SAE-SSV: Supervised Steering in Sparse Representation Spaces for Reliable Control of Language Models

### v1

@@ Abstract
Large language models (LLMs) have demonstrated impressive capabilities in natural language understanding and generation, but controlling their behavior reliably remains challenging, especially in open-ended generation settings. This paper introduces a novel supervised steering approach that operates in sparse, interpretable representation spaces. We employ sparse autoencoders (SAEs) to obtain sparse latent representations that aim to disentangle semantic attributes from model activations. Then we train linear classifiers to identify a small subspace of task-relevant dimensions in latent representations. Finally, we learn supervised steering vectors constrained to this subspace, optimized to align with target behaviors. Experiments across sentiment, truthfulness, and politics polarity steering tasks with multiple LLMs demonstrate that our supervised steering vectors achieve higher success rates with minimal degradation in generation quality compared to existing methods. Further analysis reveals that a notably small subspace is sufficient for effective steering, enabling more targeted and interpretable interventions.
@@
Zirui He1 Mingyu Jin2 Bo Shen1 Ali Payani3 Yongfeng Zhang2 Mengnan Du1 1NJIT 2Rutgers University 3Cisco
@@ 3.2 Supervised Steering Vector Optimization
where is a language modeling loss that penalizes degraded generation quality, and encourages sparsity within the steering subspace.
@@ 4.1 Experimental Setup
Baseline Methods. We compare our SAE-SSV method against four state-of-the-art steering baselines, including Concept Activation Addition (CAA) Rimsky et al. (2024), Representation Perturbation (RePe) Zou et al. (2023), Top PC Im and Li (2025), and Inference-Time Intervention (ITI) Li et al. (2024). Please refer to Section C in Appendix for more details of baselines.
@@
Steering Success Rate (SR): The percentage of generated outputs that successfully exhibit the target attribute. We use GPT-4o-mini as an automatic judge to assess whether the generated text reflects the intended attribute. Formally, , where is the number of generations judged as exhibiting the target attribute. Specific prompting details for the judge are provided in Appendix G.
@@
Implementation Details. We report main results using 16K-dimensional SAE models for both Gemma-2-2b and Gemma-2-9b models, and a 32K-dimensional SAE for LLaMA3.1-8b model. Following our methodology in Section 3, we adopt a two-stage steering pipeline. In Stage 1, we train linear probes per task to ensure stability in feature selection. We set the number of selected SAE dimensions to to ensure sufficient subspace coverage for semantic manipulation. In Stage 2, the steering vector is optimized using a contrastive objective that combines distance loss, language modeling loss, and regularization, with coefficients , , and , respectively. Please refer to Section B in Appendix for more implementation details.
@@ 4.3 Identifying a Minimal Steering Subspace
In Figure 3(b), we incrementally select the top- ranked dimensions from our identified 128-dimensional subspace and measure class separability by calculating the difference between mean projection scores of positive and negative samples. The results demonstrate that even a small number of the highest-ranked dimensions achieves substantial class separation, with diminishing returns as more dimensions are added. This suggests that within our already focused 128-dimensional concept space, an even smaller subset of dimensions carries the most significant task-relevant information. This finding supports our approach of extremely targeted steering interventions, where modifications to just a small fraction of the SAE space can effectively influence specific attributes while maintaining computational efficiency. We provide the sets of SAE features used for constructing SSVs in Table 2 and more analysis in Appendix E.
@@ 4.6 Ablation Study
Table 4 examines the impact of two key components in our method: the supervised training of the steering vector and the inclusion of the LM loss. Removing either component leads to a clear drop in steering success. Notably, omitting the LM loss increases SR to 28.6%, but also causes a substantial rise in output disorder (43.3%), indicating unstable model behavior. In contrast, the full SAE-SSV achieves the highest SR (63.2%) while maintaining low disorder (13.3%), demonstrating the importance of subspace-constrained, supervised optimization. In addition, we study the effect of the scaling factor used during inference. We observe that the steering strength measured qualitatively by semantic shift is approximately linear with respect to . However, developing a precise quantitative metric for steering intensity remains challenging. We provide representative examples illustrating this relationship in Appendix F.
@@ 5 Related Work
Steering methods for LLMs have evolved from heuristic techniques like activation addition Zou et al. (2023) to more sophisticated representation interventions. Concurrently, sparse autoencoders (SAEs) have emerged as tools for disentangling superimposed model features Bricken et al. (2023); Huben et al. (2024). Recent work has begun exploring their intersection, with Chalnev et al. (2024) targeting specific SAE features for steering, and He et al. (2025) demonstrating that SAE latents can capture instruction-following mechanisms. Feature-based steering has been applied to toxicity mitigation Gallifant et al. (2025) and safety alignment Wu et al. (2025a). For detailed related work, please refer to Section A in Appendix.
6 Conclusions
In this paper, we introduced SAE-SSV, a framework that enables effective LLM steering by operating in sparse, task-specific subspaces. The key insight lies in constraining interventions to a small number of interpretable dimensions that capture task-relevant semantics, enabling more targeted control while preserving generation quality. Experiments across sentiment, truthfulness, and political polarity steering tasks with multiple LLMs demonstrate that SAE-SSV consistently outperforms existing methods by a substantial margin.
@@ 6 Conclusions
Appendix A Related Work
Language Model Representations
Activation Steering
Activation steering has emerged as a powerful technique for influencing model behavior during inference without retraining. Early work by Dathathri et al. (2020) introduced Plug and Play Language Models, while Zou et al. (2023) systematized representation engineering across various attributes. Subsequent research has demonstrated steering effectiveness in improving truthfulness Marks and Tegmark (2024); Tigges et al. (2023), enhancing safety Arditi et al. (2024); Li et al. (2024), mitigating biases Jorgensen et al. (2024), and controlling style Wang (2024). Recent advancements include CAA Rimsky et al. (2024), which uses contrastive activation addition, RePe Kleindessner et al. (2023), employing PCA-derived directions, and ITI Li et al. (2024), which iteratively trains steering vectors. However, as noted by Zhou et al. (2024), steering often faces a trade-off between control strength and generation quality, particularly in open-ended settings. This limitation stems partly from operating in dense representation spaces, where interventions along one direction inadvertently affect many related concepts Huben et al. (2024). Our work addresses this challenge by leveraging the disentangled features learned by SAEs and employing supervised dimension selection to constrain steering to a task-specific subspace, thereby enabling more targeted interventions with minimal side effects.
Sparse Autoencoders
Sparse autoencoders address the entanglement challenge through dictionary learning, which disentangles features in superposition. By mapping representations into a higher-dimensional space with sparsity constraints, SAEs extract more interpretable features. Architectural variants include vanilla SAEs Sharkey et al. (2022) and TopK SAEs Gao et al. (2024), with pre-trained models like Gemma Scope Lieberum et al. (2024) and Llama Scope He et al. (2024) enabling broader research applications. Recent work has leveraged these SAEs to interpret model representations Kissane et al. (2024) and understand model capabilities Ferrando et al. (2025), but their application to controlled generation remains limited.
Appendix B More Implementation Details
Optimization is performed for 100 iterations using a learning rate of 0.05 and a batch size of 64. During inference, we apply the steering vector at each decoding step with scaling factors ranging from 1.0 to 10.0 to explore the trade-off between steering strength and output quality. We elaborate on parameter selection in Section 4.3. All experiments are conducted on a single NVIDIA A100 GPU.
Appendix C Baseline Methods
We compare our SAE-SSV method against several steering baselines:
Concept Activation Addition (CAA) Rimsky et al. (2024): CAA adds the mean activation difference between positive and negative examples during inference to steer model outputs.
Representation Perturbation (RePe) Zou et al. (2023): It perturbs activations along principal components of class-conditional differences.
Top PC Im and Li (2025): This method projects activations onto the first principal component of the embedding space, capturing the direction of maximal variance.
Inference-Time Intervention (ITI) Li et al. (2024): This steering method shifts attention head activations during inference along truth-related directions found via linear probing.
Appendix D Case Study
@@
Appendix E SAE-SSV Features Analysis
@@
Appendix F Intervention Factors
@@
Appendix G Evalutation Method Details

### v2

@@ Abstract
Large language models (LLMs) have demonstrated impressive capabilities in natural language understanding and generation, but controlling their behavior reliably remains challenging, especially in open-ended generation settings. This paper introduces a novel supervised steering approach that operates in sparse, interpretable representation spaces. We employ sparse autoencoders (SAEs) to obtain sparse latent representations that aim to disentangle semantic attributes from model activations. Then we train linear classifiers to identify a small subspace of task-relevant dimensions in latent representations. Finally, we learn supervised steering vectors constrained to this subspace, optimized to align with target behaviors. Experiments across sentiment, truthfulness, and politics polarity steering tasks with multiple LLMs demonstrate that our supervised steering vectors achieve higher success rates with minimal degradation in generation quality compared to existing methods. Further analysis reveals that a notably small subspace is sufficient for effective steering, enabling more targeted and interpretable interventions. Our implementation is publicly available at https://github.com/Ineedanamehere/SAE-SSV.
@@
Zirui He1 Mingyu Jin2 Bo Shen1 Ali Payani3 Yongfeng Zhang2 Mengnan Du1††thanks: Corresponding author. 1NJIT 2Rutgers University 3Cisco
@@ 3.2 Supervised Steering Vector Optimization
where is a language modeling loss that penalizes degraded generation quality by computing the cross-entropy of the positive target sequence conditioned on the steered hidden state of the negative input , and encourages sparsity within the steering subspace.
@@ 4.1 Experimental Setup
Baseline Methods. We compare our SAE-SSV method against four widely used steering baselines:
Concept Activation Addition (CAA) Rimsky et al. (2024): Adds the mean activation difference between positive and negative examples during inference to steer model outputs.
Representation Perturbation (RePe) Zou et al. (2023): Perturbs activations along principal components of class-conditional differences.
Top PC Im and Li (2025): Projects activations onto the first principal component of the embedding space, capturing the direction of maximal variance.
Inference-Time Intervention (ITI) Li et al. (2024): Shifts attention head activations during inference along truth-related directions found via linear probing.
@@
Steering Success Rate (SR): The percentage of generated outputs that successfully exhibit the target attribute. We use GPT-4o-mini as an automatic judge to assess whether the generated text reflects the intended attribute. Formally, , where is the number of generations judged as exhibiting the target attribute. Specific prompting details for the judge are provided in Appendix D.
@@
Implementation Details. We report main results using 16K-dimensional SAE models for both Gemma-2-2b and Gemma-2-9b models, and a 32K-dimensional SAE for LLaMA3.1-8b model. Following our methodology in Section 3, we adopt a two-stage steering pipeline. In Stage 1, we train linear probes per task to ensure stability in feature selection. We set the number of selected SAE dimensions to to ensure sufficient subspace coverage for semantic manipulation. In Stage 2, the steering vector is optimized using a contrastive objective that combines distance loss, language modeling loss, and regularization, with coefficients , , and , respectively. Optimization is performed for 100 iterations with a learning rate of 0.05 and a batch size of 64. During inference, we apply the steering vector at each decoding step with scaling factors ranging from 1.0 to 10.0 to explore the trade-off between steering strength and output quality. For each model and task, we apply interventions at empirically selected layers: LLaMA3.1-8B (layer 16 for all tasks), Gemma2-2B (sentiment: 13, truthfulness: 16, politics: 15), and Gemma2-9B (sentiment: 20, truthfulness: 26, politics: 20). All experiments are conducted on a single NVIDIA A100 GPU.
@@ 4.3 Identifying a Minimal Steering Subspace
In Figure 3(b), we incrementally select the top- ranked dimensions from our identified 128-dimensional subspace and measure class separability by calculating the difference between mean projection scores of positive and negative samples. The results demonstrate that even a small number of the highest-ranked dimensions achieves substantial class separation, with diminishing returns as more dimensions are added. This suggests that within our already focused 128-dimensional concept space, an even smaller subset of dimensions carries the most significant task-relevant information. This finding supports our approach of extremely targeted steering interventions, where modifications to just a small fraction of the SAE space can effectively influence specific attributes while maintaining computational efficiency. We provide the sets of SAE features used for constructing SSVs in Table 2 and more analysis in Appendix B.
@@ 4.6 Ablation Study
Table 4 examines the impact of two key components in our method: the supervised training of the steering vector and the inclusion of the LM loss. Removing either component leads to a clear drop in steering success. Notably, omitting the LM loss increases SR to 28.6%, but also causes a substantial rise in output disorder (43.3%), indicating unstable model behavior. In contrast, the full SAE-SSV achieves the highest SR (63.2%) while maintaining low disorder (13.3%), demonstrating the importance of subspace-constrained, supervised optimization. In addition, we study the effect of the scaling factor used during inference. We observe that the steering strength measured qualitatively by semantic shift is approximately linear with respect to . However, developing a precise quantitative metric for steering intensity remains challenging. We provide representative examples illustrating this relationship in Appendix C.
@@ 5 Related Work
Language Model Representations.
Activation Steering.
Activation steering has emerged as a powerful technique for influencing model behavior during inference without retraining. Early work such as Plug and Play Language Models Dathathri et al. (2020) and representation engineering Zou et al. (2023) established the feasibility of direct activation manipulation. Subsequent research demonstrated its effectiveness in improving truthfulness Marks and Tegmark (2024); Tigges et al. (2023), enhancing safety Arditi et al. (2024); Li et al. (2024), mitigating biases Jorgensen et al. (2024), and controlling style Wang (2024). More recent methods include CAA Rimsky et al. (2024), which uses contrastive activation addition, RePe Kleindessner et al. (2023), which employs PCA-derived directions, and ITI Li et al. (2024), which iteratively trains steering vectors. Nevertheless, steering often faces a trade-off between control strength and generation quality in open-ended settings Zhou et al. (2024), in part because interventions in dense spaces can inadvertently entangle multiple concepts Huben et al. (2024). Our work addresses this challenge by leveraging disentangled SAE features and supervised dimension selection to constrain steering to a task-specific subspace, enabling more targeted interventions with fewer side effects.
Sparse Autoencoders.
Sparse autoencoders (SAEs) have been introduced to disentangle superimposed features through dictionary learning. By mapping activations into a higher-dimensional sparse space, SAEs yield more interpretable features Bricken et al. (2023); Huben et al. (2024). Variants include vanilla SAEs Sharkey et al. (2022) and TopK SAEs Gao et al. (2024), with pre-trained repositories such as Gemma Scope Lieberum et al. (2024) and Llama Scope He et al. (2024) enabling broader research. SAEs have been used to interpret model representations Kissane et al. (2024), to understand model capabilities Ferrando et al. (2025), and to explore intersections with steering Chalnev et al. (2024); He et al. (2025). Applications include toxicity mitigation Gallifant et al. (2025) and safety alignment Wu et al. (2025a), but the use of SAEs for controllable generation remains relatively limited. Our work extends this line by combining SAE-derived features with supervised optimization to construct effective steering vectors.
6 Conclusions and Future Work
In this paper, we introduced SAE-SSV, a framework that enables effective LLM steering by operating in sparse, task-specific subspaces. The key insight lies in constraining interventions to a small number of interpretable dimensions that capture task-relevant semantics, enabling more targeted control while preserving generation quality. Experiments across sentiment, truthfulness, and political polarity steering tasks with multiple LLMs demonstrate that SAE-SSV consistently outperforms existing methods by a substantial margin. Our cross dataset experiments reveal that SAE-SSV captures both semantic directions and stylistic patterns of training data, highlighting its potential as a more general steering mechanism. For our future work, we aim to achieve universal and style-invariant SSVs that generalize across datasets, tasks, and model families by curating diverse training corpora and developing objectives that explicitly encourage semantic steering while minimizing sensitivity to stylistic variation.
@@ 6 Conclusions and Future Work
Acknowledgments
Mengnan Du is supported by National Science Foundation (NSF) Grant #2310261. The views and conclusions in this paper are those of the authors and should not be interpreted as representing any funding agencies.
Appendix A Case Study
@@
Appendix B SAE-SSV Features Analysis
@@
Appendix C Intervention Factors
@@
Appendix D Evalutation Method Details